The local community college theater put on a production of chicago. There were 186 tickets sold, some for $16 ( non students price ) and others for $12 (students price ) . If the receipts for one performance totaled $2640, How much of each type of ticket were sold?


1
Expert's answer
2021-10-06T00:36:23-0400

If x is the quantity sold at non-student price and y is quantity sold at student price, then we can make the system of two equations:

x + y = 186,

16x + 12y = 2640,

x = 186 - y,

16×(186 - y) + 12y = 2640,

2976 - 4y = 2640,

4y = 336,

y = 84 tickets sold at student price,

x = 186 - 84 = 102 tickets sold at non-student price.